OpenStudy (anonymous):

can x^2+x+2 be factored?

OpenStudy (anonymous):

Not with real numbers.

OpenStudy (anonymous):

\[{-1 \pm \sqrt{1^2 -(4 \times 1 \times 2)} \over 2} = {-1 \pm \sqrt{-7} \over 2} = {-1\pm \sqrt{7i} \over 2} \]

OpenStudy (apoorvk):

Nope. To check if a quadratic has real factors or not, check the Discriminant "D" \[D = b^2 - 4ac\] where a,b and ccan be found by comparing your quadratic with the standard form: \[ax^2 + bx +c = 0\] now, if: D=0. --->One real root or factor (that means its a complete square) D>0 --->Two real factors D<0 ---> NO real factors So, what do you think now for the above problem?

OpenStudy (anonymous):

I still say it does ;)

OpenStudy (apoorvk):

It doesn't. Not real factors atleast. :)

OpenStudy (anonymous):

how do I solve this problem?

OpenStudy (apoorvk):

Cross multiply. best way.

OpenStudy (anonymous):

so would the answer be x^2 over c+1?

OpenStudy (anonymous):

the answer given is 1... i have no clue how